OpenStudy (anonymous):

What is the equation of a line that passes through (7, 8) and has a slope of -3?

OpenStudy (anonymous):

help please

OpenStudy (sleepyhead314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) so plug in -3 for m 7 for x1 and 8 for y1

OpenStudy (anonymous):

thanks :) it gave me my answer

OpenStudy (sleepyhead314):

you're welcome :)